OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2

OpenStudy (ivettef365):

where are the equations

OpenStudy (anonymous):

y - 4 = (x - 1) y + 4 = 2(x - 1) y + 4 = (x - 1) y - 4 = 2(x + 1)

OpenStudy (ivettef365):

are you sure the slope = 1/2 ?

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

formula is y - y1 = m (x - x1)

OpenStudy (ivettef365):

you have (1,-4) with slope of 1/2 should be y + 4 = 1/2(x-1)

OpenStudy (ivettef365):

maybe the choices are wrong ?

OpenStudy (anonymous):

oh im sorry i forgot to include the 1/2s in the equations but the answer you have is correct so thanks

OpenStudy (ivettef365):

yw :)